OpenStudy (anonymous):

I just want to check my answer.. The gradient of a curve is given by dy/dx = kx^2+6, where k is a constant, given that the curve passes through (1,6) and (-1,2). Find the value of y in terms of x. I have found k= 15 so it's y= 15x^3/3 + 6x +c or y= 5x^3 + 6x + c .

OpenStudy (anonymous):

by integration.

OpenStudy (anonymous):

i got k=-12 and c=4. Check your work again. I know mine's right cuz you can easily verify it by plugging the values into the equation.

OpenStudy (anonymous):

oh okay , imma check it

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

wrong answer :/ I did like this : integral (kx^2+6) dx y = kx^2/3 + 6x + c 6 = k(1)^2/3 + 6 (1) + c ---> (1) 2 = k(-1)^2/3 + 6 (-1) + c ---> (2) by subtracting (1) - (2) 2 = 2/3k + 12 2-12 = 2/3k k= -15 .. ?

OpenStudy (anonymous):

and c=3

OpenStudy (anonymous):

what's the integral of x^3?

OpenStudy (anonymous):

x^4/4 ?

OpenStudy (anonymous):

sorry. the integral of x^2?

OpenStudy (anonymous):

x^3/3 .. I just noticed that^ But I calculated it x^3/3

OpenStudy (anonymous):

6-2?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

yeah but that's the mistake. you've used 2 in your calculation...

OpenStudy (anonymous):

oh pellet , silly mistakes can flutter the whole calculation :s

OpenStudy (anonymous):

yeah lol

OpenStudy (anonymous):

sorry I didn't concentrate, thanks :D

OpenStudy (anonymous):

word filters - you gotta hate them ;)

OpenStudy (anonymous):

haha yeah F U C K IT ;)

OpenStudy (anonymous):

lol